# M. C. Rodriguez

### v1

@@ 4 Case 2
Our results are shown in Figs.(2,3,4,5) where we can say
@@
However if we take $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ not no null the Majoron dissapear but it has mass of few GeV, the same results as presented in [21], but the values of $M_{A}$ is near the ones we present in our Figs.(2,4).
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric 3-3-1 model with all the VEV from the sextet and anti-sextet not no null and we shown our light scalar fields has mass near the value $M_{h}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(21), and $M_{A}>150\,\ {\mbox{G}eV}$.
@@ Acknowledgments
We would like thanks V. Pleitez for useful discussions above 3-3-1 models. We also to thanks IFT for the nice hospitality during my several visit to perform my studies about the severals 3-3-1 Models and also for done this article.

### v3

@@ 4 Case 2
Our results are shown in Figs.(7,8,9,10) where we can say
@@
However if we take $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ not no null the Majoron dissapear but it has mass of few GeV, the same results as presented in [25], but the values of $M_{A}$ is near the ones we present in our Figs.(7,9).
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet not no vanishing and we shown our light scalar fields has mass near the value $M_{h}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(15), and $M_{A}>150\,\ {\mbox{G}eV}$.
@@ Acknowledgments
We would like thanks V. Pleitez for useful discussions above $331$ models. We also to thanks IFT for the nice hospitality during my several visit to perform my studies about the severals $331$ Models and also for done this article.